# The Poisson Midpoint Method for Langevin Dynamics: Provably Efficient Discretization for Diffusion Models

### v2

@@ 3 Main Results
Let $X_{t}$ be the iterates of $\mathsf{S}(A,G,\Gamma,b,\tfrac{\alpha}{K})$ and $X^{\mathsf{P}}_{t}$ be the iterates of $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ with OPTION 1. Suppose that $X_{0}^{\mathsf{P}}=X_{0}$. Let $\tilde{X}_{tK+i}$ be the iterates in Equation (8). Define random variables:
@@
$\quad\beta_{tK+i}:=\|K\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i},t\alpha+\tfrac{\alpha i}{K})-b(\tilde{X}_{tK},t\alpha)]\|$
@@
$\leq\sum_{s=0}^{T-1}\sum_{i=0}^{K-1}\mathbb{E}[\|B_{sK+i}\|^{2}]+C\mathbb{E}\left[\frac{\beta^{4}_{sK+i}}{K^{2}}+\frac{\beta^{10}_{sK+i}}{K^{3}}+\frac{\beta^{6}_{sK+i}}{K^{2}}+\frac{\beta^{2r}_{sK+i}}{K}\right]$ (10)
@@
$F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L$-smooth (i.e., $\nabla F$ is $L$-Lipschitz). $\mathbf{x}^{*}$ is its global minimizer.
@@
The assumptions above are very mild and standard in the literature. Specifically, Assumption 2 shows that the initialization is close to global optimum by $O(\sqrt{d})$ up to 14th moments. For instance, this is satisfied when the initialization is a standard Gaussian variable with mean $\mu$ satisfying $\|\mu-\mathbf{x}^{*}\|=O(\sqrt{d})$. Specifically this is true when $\mu=0$ and $\|\mathbf{x}^{*}\|=O(\sqrt{d})$. This is a weak assumption which is implied from common initialization assumptions in the literature as listed below. It can be replaced with the assumptions in [53, Appendix D and Lemma 27] which considers Gaussian initializations with the right variance and mean. The original randomized midpoint method work [40] considers initializing at $\mathbf{x}^{*}$ whereas [45] considers a Gaussian initialization with the right variance at a local minimum of $F$.
We do not make any assumptions regarding isoperimetry of the target distribution $\pi^{\star}(\mathbf{x})\propto\exp(-F(\mathbf{x}))$.
@@
$\quad+O(CL^{4}\alpha^{4}Kd^{2}T)$ (11)
Remark 2.
There are lower order terms hidden in the $O()$ notation. These are explicated in Equation (40) in the appendix. The next theorem gives a similar guarantee for ULMC and the lower order terms are explicated in Equation 60 in the appendix.
@@
$\quad+O(\tfrac{K\alpha^{7}L^{4}T^{2}}{\gamma}(d+\log K)^{2})$ (12)
@@
Table 1: Comparison of LMC and PLMC guarantees. LMC complexity is the upper bound on the number of drift ($b()$) evaluations to achieve the error guarantee in the referenced work. PLMC complexity is the corresponding upper bound for PLMC. PLMC obtains a quadratic improvement in $\epsilon$, and improved dependence on $\frac{L}{\lambda},d$. The bounds hold up to poly-log factors.
Algorithm Reference Conditions LMC complexity PLMC complexity
@@ 4 Proof Sketch
Where $\tilde{Z}_{tK+i}:=Z_{tK+i}+B_{tK+i}+S_{tK+i}$, $B_{tK+i}$ is the ‘bias’ with a non-zero conditional mean, and $S_{tK+i}$ is the variance with 00 conditional mean (conditioned on $\tilde{X}_{tK+i}$). They are independent of $Z_{tK+i}$ conditioned on $\tilde{X}_{tK+i}$. Note that the sequence $(\tilde{Z}_{tK+i})_{t,i}$ is neither i.i.d. nor Gaussian. If it was a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$, then this is exactly same as $\mathsf{S}(A,G,\Gamma,\tfrac{\alpha}{K})$.
@@
$\mathbb{E}F(\tilde{X}_{(t+1)K})-\mathbb{E}F(\tilde{X}_{tK})\lesssim-\alpha\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d+o(\alpha d)$
$\implies\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim\frac{F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x})}{\alpha}+LTd+o(LTd)$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\lesssim L^{p-1}\mathbb{E}\frac{(F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x}))^{p}}{\alpha}+TL^{p}d^{p}(1+(\alpha LT)^{p-1})$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\tilde{V}_{tK}\|^{2p}$ $\lesssim\tfrac{1}{\gamma\alpha}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2p}+\mathbb{E}|(F(\Psi_{0})-F(\mathbf{x}^{*})|^{p}\right]+T\left[\tfrac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right]d^{p}$
@@ 5 Experiments
We now present experiments to evaluate Poisson Midpoint Method as a training-free scheduler for diffusion models. We consider the Latent Diffusion Model (LDM) [39] for CelebAHQ 256, LSUN Churches, LSUN Bedrooms and FFHQ datasets using the official (PyTorch) codebase and checkpoints. We compare the sample quality of the Poisson Midpoint Method against established methods such as DDPM, DDIM and DPM-Solver, varying the number of neural network calls (corresponding to the drift $b(\mathbf{x},t)$) used to generate a single image.
To evaluate the quality, we generate $50$k images for each method and number of neural network calls and compare it with the training dataset. We use Fréchet Inception Distance (FID) [19] metric for LSUN Churches and LSUN Bedrooms. For CelebAHQ 256 and FFHQ, we use Clip-FID, a more suitable metric as it is known that FID may exhibit inconsistencies with human evaluations datasets outside of Imagenet [24]. We refer to Section A.3 in the appendix for further details.
[…]
@@ 6 Conclusion
We introduce the Poisson Midpoint Method, which efficiently discretizes Langevin Dynamics and theoretically demonstrates quadratic speed up over Euler-Maruyama discretization under general conditions. We apply our method to diffusion models for image generation, and show that our method maintains the quality of 1000 step DDPM with just 50-80 neural network calls. This outperforms ODE based methods such as DPM-Solver in terms of quality, with a similar amount of compute. Future work can explore variants of Poisson midpoint method with better performance when fewer than 50 neural network calls are used. An interesting theoretical direction would be to derive convergence bounds for algorithms such as DDPM which have a time dependent drift function. Future work can also consider convergence rates of PLMC under conditions such as the Poincare Inequality and whenever $\nabla F$ is Hölder continuous instead of Lipschitz continuous.
@@ 7 Societal Impact
We thank Prateek Jain for helpful comments during the course of this research work. This work was done when SK was a student researcher at Google Research.
@@ A.1 Pseudocode
    We take the dynamics to go backward in time (i.e., compute $X_{t-1}$ from $X_{t}$)
@@
$X_{t-1}=a_{t}X_{t}+b_{t}b(X_{t},t)+\sigma_{t}Z_{t};\quad Z_{t}\sim\mathcal{N}(0,\mathbf{I})$
@@ Appendix C Proof of Theorem 1
The proof relies on the Wasserstein CLT based approach introduced in [10] to compare two discrete time stochastic processes. For the sake of clarity we consider the drift function $b(x,t)$ to be $b(x)$ (i.e., time invariant). However, the proof goes through for time varying drifts as well.
@@ C.1 The Bias-Variance Decomposition:
$S_{tK+i}:=K(H_{t,i}-\tfrac{1}{K})\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})]$
@@ C.2 Random Function Representation
$X_{tK+i+1}=A_{\tfrac{\alpha}{K}}X_{tK+i}+G_{\tfrac{\alpha}{K}}\left[b(X_{tK+i})\right]+\Gamma_{\tfrac{\alpha}{K}}Z_{tK+i}$ (14)
@@
$\tilde{X}_{tK+i+1}=A_{\tfrac{\alpha}{K}}\tilde{X}_{tK+i}+G_{\tfrac{\alpha}{K}}\left[b(\tilde{X}_{tK+i})\right]+\Gamma_{\tfrac{\alpha}{K}}\tilde{Z}_{tK+i}$ (15)
@@
$\mathsf{KL}\left(p\boldsymbol{\bigr{|}\bigr{|}}q\right)=\mathsf{KL}\left(p_{x}\boldsymbol{\bigr{|}\bigr{|}}q_{x}\right)+\mathbb{E}_{x\sim p_{x}}\mathsf{KL}\left(p_{y|x}\boldsymbol{\bigr{|}\bigr{|}}q_{y|x}\right)$
@@
$\mathsf{KL}\left(F_{\#}P\boldsymbol{\bigr{|}\bigr{|}}F_{\#}Q\right)\leq\mathsf{KL}\left(P\boldsymbol{\bigr{|}\bigr{|}}Q\right)$
@@
$(\tilde{X}_{\tau})_{0\leq\tau\leq T}=F_{T}(X_{0},\tilde{Z}_{0},\dots,\tilde{Z}_{T-1})$
@@
$\quad\sum_{\tau=1}^{T-1}\mathbb{E}_{(X_{0},\tilde{Z}_{0:\tau-1})}\mathsf{KL}\left(\mathsf{Law}(\tilde{Z_{\tau}}|\tilde{Z}_{0:\tau-1},X_{0})\boldsymbol{\bigr{|}\bigr{|}}\mathsf{Law}(Z_{\tau})\right)$ (16)
@@ C.3 Controlling KL Divergence via Wasserstein Distances
$\mathcal{W}_{2}(\mathsf{Law}(\mathbf{A}),\mathsf{Law}(\mathbf{B}+\mathbf{C}))\leq\mathcal{W}_{2}(\mathsf{Law}(\mathbf{A}),\mathsf{Law}(\mathbf{B}))+\sqrt{\mathbb{E}\|\mathbf{C}\|^{2}}$
@@
$\leq\mathcal{W}_{2}^{2}\left(\mathsf{Law}(Z_{\tau,2}),\mathsf{Law}(Z_{\tau,2}+\sqrt{2}S_{\tau}|\tilde{Z}_{0:\tau-1},X_{0})\right)+2\mathbb{E}[\|B_{\tau}\|^{2}|\tilde{Z}_{0:\tau-1},X_{0}]$ (17)
@@ C.4 Bounding the Error Along the Trajectory:
$\left(\mathcal{W}_{2}\left(P,Q\right)\right)^{2}$ $\leq 3(1+\nu)\left[\beta^{4}\nu^{3}+\beta^{2}\nu^{2}\right]\exp(\tfrac{3\beta^{2}}{2})$ (18)
@@
$\left(\mathcal{W}_{2}\left(P,Q\right)\right)^{2}\leq 2\nu$ (19)
@@
[…]
@@
$\mathcal{W}_{2}^{2}(\mathcal{N}(0,\Sigma_{1}),\mathcal{N}(0,\Sigma_{2}))=\mathsf{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{\tfrac{1}{2}}\Sigma_{1}\Sigma_{2}^{\tfrac{1}{2}})^{\tfrac{1}{2}})$
@@
$\bar{Q}:=\mathsf{Law}(Z_{Kt+i,2}+\sqrt{2}(S_{Kt+i})|X_{0},\tilde{Z}_{0:Kt+i-1})$
@@
$\mathcal{W}_{2}^{2}(\bar{P},\bar{Q})\leq C\mathbb{E}\left[\tfrac{\beta^{4}}{K^{2}}+\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}+\tfrac{\beta^{2r}}{K}\bigr{|}\tilde{Z}_{0:Kt+i-1},X_{0}\right]$ (20)
@@
$\mathbb{E}[\sqrt{2}S_{tK+i}|\tilde{Z}_{0:Kt+i-1},Y_{0:Kt+i-1},X_{0}]=0$
@@
$\beta:=\|K\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})]\|$ (21)
@@
$\Sigma:=2\mathbb{E}[S_{tK+i}S_{tK+i}^{\intercal}|\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1}X_{0}]$ (22)
@@
$\sqrt{2}S_{tK+i}$ takes its values in a one-dimensional sub-space almost surely when conditioned on $\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1},X_{0}$ It is independent of $Z_{tK+i}$ conditioned on $\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1},X_{0}$.
Let us now define the following random probability distributions measurable with respect to the sigma algebra of $\tilde{Z}_{0:Kt+i-1},X_{0}$
@@
    $Q:=\mathsf{Law}(Z_{Kt+i,2}+\sqrt{2}S_{Kt+i}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$
@@
    $P:=\mathsf{Law}(\sqrt{I+\Sigma}Z_{Kt+i,2}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$
@@
$\mathcal{W}_{2}^{2}\left(\bar{P},\bar{Q}\right)\leq\mathbb{E}[\mathcal{W}_{2}^{2}\left(\bar{P},Q\right)|\tilde{Z}_{0:Kt+i-1},X_{0}]$
$\leq 2\mathbb{E}[\mathcal{W}_{2}^{2}\left(\bar{P},P\right)+\mathcal{W}_{2}^{2}\left(P,Q\right)|\tilde{Z}_{0:Kt+i-1},X_{0}]$ (23)
@@
$\leq C\left[\left[\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}\right]\mathbbm{1}(\beta\leq 1)+\frac{\beta^{2}}{K}\mathbbm{1}(\beta>1)\right]$ (25)
@@
$\leq C\left[\tfrac{\beta^{4}}{K^{2}}+\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}+\tfrac{\beta^{2r}}{K}\right]$ (26)

### v3

@@ 3 Main Results
Let $X_{t}$ be the iterates of $\mathsf{S}(A,G,\Gamma,\tfrac{\alpha}{K})$ and $X^{\mathsf{P}}_{t}$ be the iterates of $\mathsf{PS}(A,G,\Gamma,\alpha,K)$ with OPTION 1. Suppose that $X_{0}^{\mathsf{P}}=X_{0}$. Let $\tilde{X}_{tK+i}$ be the iterates in Equation (8). Define random variables:
@@
$\quad\beta_{tK+i}:=\|K\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i},t\alpha+\tfrac{\alpha i}{K})-b(\tilde{X}_{tK},t\alpha)]\|.$
@@
$\leq\sum_{s=0}^{T-1}\sum_{i=0}^{K-1}\mathbb{E}[\|B_{sK+i}\|^{2}]+C\mathbb{E}\left[\frac{\beta^{4}_{sK+i}}{K^{2}}+\frac{\beta^{10}_{sK+i}}{K^{3}}+\frac{\beta^{6}_{sK+i}}{K^{2}}+\frac{\beta^{2r}_{sK+i}}{K}\right].$ (10)
@@
$F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L$-smooth (i.e, $\nabla F$ is $L$-Lipschitz). $\mathbf{x}^{*}$ is its global minimizer.
@@
The assumptions above are very mild and standard in the literature. We do not make any assumptions regarding isoperimetry of the target distribution $\pi^{\star}(\mathbf{x})\propto\exp(-F(\mathbf{x}))$.
@@
$\quad+CL^{4}\alpha^{4}Kd^{2}T+\mathrm{Lower\ order\ terms}.$ (11)
The lower order terms are explicated in Equation (40). The next theorem gives a similar guarantee for ULMC and the lower order terms are explicated in Equation 60.
@@
$\quad+\tfrac{K\alpha^{7}L^{4}T^{2}}{\gamma}(d+\log K)^{2}+\mathrm{Lower\ order\ terms}.$ (12)
@@
Table 1: Comparison of LMC and PLMCguarantees. LMC steps is the upper bound for $T$ for LMC to achieve the error guarantee in the referenced work. PLMC steps is the corresponding upper bound for PLMC. PLMC obtains a quadratic improvement in $\epsilon$, and improved dependence on $\frac{L}{\lambda},d$.
Algorithm Reference Conditions LMC steps PLMC steps (ours)
@@ 4 Proof Sketch
$\tilde{X}_{tK+i+1}=A_{\tfrac{\alpha}{K}}\tilde{X}_{tK+i}+G_{\tfrac{\alpha}{K}}\left[b(\tilde{X}_{tK+i})\right]+\Gamma_{\tfrac{\alpha}{K}}\tilde{Z}_{tK+i}.$
Here $\tilde{Z}_{tK+i}:=Z_{tK+i}+B_{tK+i}+S_{tK+i}$, $B_{tK+i}$ is the ‘bias’ with a non-zero conditional mean, and $S_{tK+i}$ is the variance with 0 conditional mean (conditioned on $\tilde{X}_{tK+i}$). They are independent of $Z_{tK+i}$ conditioned on $\tilde{X}_{tK+i}$. Note that the sequence $(\tilde{Z}_{tK+i})_{t,i}$ is neither i.i.d. nor Gaussian. If it was a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$, then this is exactly same as $\mathsf{S}(A,G,\Gamma,\tfrac{\alpha}{K})$.
@@
$\mathbb{E}F(\tilde{X}_{(t+1)K})-\mathbb{E}F(\tilde{X}_{tK})\lesssim-\alpha\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d+\text{ Lower order terms}$
$\implies\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim\frac{F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x})}{\alpha}+LTd+\text{ Lower order terms}.$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\lesssim L^{p-1}\mathbb{E}\frac{(F(X_{0})-\inf_{\mathbf{x}}F(\mathbf{x}))^{p}}{\alpha}+TL^{p}d^{p}(1+(\alpha LT)^{p-1})+\text{Lower order terms}.$
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\tilde{V}_{tK}\|^{2p}$ $\lesssim\tfrac{1}{\gamma\alpha}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2p}+\mathbb{E}|(F(\Psi_{0})-F(\mathbf{x}^{*})|^{p}\right]+T\left[\tfrac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right]d^{p}.$
@@ 5 Experiments
We now present experiments to evaluate Poisson Midpoint Method as a training-free scheduler for diffusion models. We consider the Latent Diffusion Model (LDM) [39] for CelebAHQ 256, LSUN Churches, LSUN Bedrooms and FFHQ datasets using the official (PyTorch) codebase and checkpoints. We compare the sample quality of the Poisson Midpoint Method against established methods such as DDPM, DDIM and DPMSolver, varying the number of neural network calls (corresponding to the drift $b(\mathbf{x},t)$) used to generate a single image.
To evaluate the quality, we generate 5050k images for each method and number of neural network calls and compare it with the training dataset. We use Fréchet Inception Distance (FID) [19] metric for LSUN Churches and LSUN Bedrooms. For CelebAHQ 256 and FFHQ, we use Clip-FID, a more suitable metric as it is known that FID may exhibit inconsistencies with human evaluations datasets outside of Imagenet [24]. We refer to Section A.3 in the appendix for further details.
We refer to the ODE based sampler with $\eta=0$ (see [42]) setting as DDIM and use the implementation in [39]. We generate images for number of neural network calls ranging from 20 to 500. For DPMSolver, we port the official codebase of [28] to generate images for different numbers of neural network calls ranging from 10 to 100 using MultistepDPMSolver and tune the hyperparameter ‘order’ over $\{2,3\}$ and ‘skip_type’ over $\{\text{`logSNR'},\text{`time\_uniform'},\text{`time\_quadratic'}\}$ for each instance to obtain the best possible FID score. This ensures that the baseline is competitive.
@@ 6 Conclusion
We introduce the Poisson Midpoint Method, which efficiently discretizes Langevin Dynamics and theoretically demonstrates quadratic speed up over Euler-Maruyama discretization under general conditions. We apply our method to diffusion models for image generation, and show that our method maintains the quality of 1000 step DDPM with just 50-80 neural network calls. This outperforms ODE based methods such as DPM-Solver in terms of quality, with a similar amount of compute. Future work can explore variants of Poisson midpoint method with better performance when fewer than 50 neural network calls are used.
@@ 7 Societal Impact
We thank Prateek Jain for helpful comments during the course of this research work.
@@ A.1 Pseudocode
    We take the dynamics to go backward in time (i.e, compute $X_{t-1}$ from $X_{t}$).
@@
$X_{t-1}=a_{t}X_{t}+b_{t}b(X_{t},t)+\sigma_{t}Z_{t};\quad Z_{t}\sim\mathcal{N}(0,\mathbf{I}).$
@@ Appendix C Proof of Theorem 1
The proof relies on the Wasserstein CLT based approach introduced in [10] to compare two discrete time stochastic processes. For the sake of clarity we consider the drift function $b(x,t)$ to be $b(x)$ (i.e, time invariant). However, the proof goes through for time varying drifts as well.
@@ C.1 The Bias-Variance Decomposition:
$S_{tK+i}:=K(H_{t,i}-\tfrac{1}{K})\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})].$
@@ C.2 Random Function Representation
$X_{tK+i+1}=A_{\tfrac{\alpha}{K}}X_{tK+i}+G_{\tfrac{\alpha}{K}}\left[b(X_{tK+i})\right]+\Gamma_{\tfrac{\alpha}{K}}Z_{tK+i}.$ (14)
@@
$\tilde{X}_{tK+i+1}=A_{\tfrac{\alpha}{K}}\tilde{X}_{tK+i}+G_{\tfrac{\alpha}{K}}\left[b(\tilde{X}_{tK+i})\right]+\Gamma_{\tfrac{\alpha}{K}}\tilde{Z}_{tK+i}.$ (15)
@@
$\mathsf{KL}\left(p\boldsymbol{\bigr|\bigr|}q\right)=\mathsf{KL}\left(p_{x}\boldsymbol{\bigr|\bigr|}q_{x}\right)+\mathbb{E}_{x\sim p_{x}}\mathsf{KL}\left(p_{y|x}\boldsymbol{\bigr|\bigr|}q_{y|x}\right).$
@@
$\mathsf{KL}\left(F_{\#}P\boldsymbol{\bigr|\bigr|}F_{\#}Q\right)\leq\mathsf{KL}\left(P\boldsymbol{\bigr|\bigr|}Q\right).$
@@
$(\tilde{X}_{\tau})_{0\leq\tau\leq T}=F_{T}(X_{0},\tilde{Z}_{0},\dots,\tilde{Z}_{T-1}).$
@@
$\quad\sum_{\tau=1}^{T-1}\mathbb{E}_{(X_{0},\tilde{Z}_{0:\tau-1})}\mathsf{KL}\left(\mathsf{Law}(\tilde{Z_{\tau}}|\tilde{Z}_{0:\tau-1},X_{0})\boldsymbol{\bigr|\bigr|}\mathsf{Law}(Z_{\tau})\right).$ (16)
@@ C.3 Controlling KL Divergence via Wasserstein Distances
$\mathcal{W}_{2}(\mathsf{Law}(\mathbf{A}),\mathsf{Law}(\mathbf{B}+\mathbf{C}))\leq\mathcal{W}_{2}(\mathsf{Law}(\mathbf{A}),\mathsf{Law}(\mathbf{B}))+\sqrt{\mathbb{E}\|\mathbf{C}\|^{2}}.$
@@
$\leq\mathbb{E}\mathcal{W}_{2}^{2}\left(\mathsf{Law}(Z_{\tau,2}),\mathsf{Law}(Z_{\tau,2}+\sqrt{2}S_{\tau}|\tilde{Z}_{0:\tau-1},X_{0})\right)+\mathbb{E}[\|B_{\tau}\|^{2}|\tilde{Z}_{0:\tau-1},X_{0}].$ (17)
@@ C.4 Bounding the Error Along the Trajectory:
$\left(\mathcal{W}_{2}\left(P,Q\right)\right)^{2}$ $\leq 3(1+\nu)\left[\beta^{4}\nu^{3}+\beta^{2}\nu^{2}\right]\exp(\tfrac{3\beta^{2}}{2}).$ (18)
@@
$\left(\mathcal{W}_{2}\left(P,Q\right)\right)^{2}\leq 2\nu.$ (19)
@@
$\mathcal{W}_{2}\left(\mu,\int Q(\cdot,\omega)dM(\omega)\right)\leq\int\mathcal{W}_{2}\left(\mu,Q(\cdot,\omega)\right)dM(\omega).$
@@
$\mathcal{W}_{2}^{2}(\mathcal{N}(0,\Sigma_{1}),\mathcal{N}(0,\Sigma_{2}))=\mathsf{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{\tfrac{1}{2}}\Sigma_{1}\Sigma_{2}^{\tfrac{1}{2}})^{\tfrac{1}{2}}).$
@@
$\bar{Q}:=\mathsf{Law}(Z_{Kt+i,2}+\sqrt{2}(S_{Kt+i})|X_{0},\tilde{Z}_{0:Kt+i-1}).$
@@
$\mathcal{W}_{2}^{2}(\bar{P},\bar{Q})\leq C\mathbb{E}\left[\tfrac{\beta^{4}}{K^{2}}+\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}+\tfrac{\beta^{2r}}{K}\bigr|\tilde{Z}_{0:Kt+i-1},X_{0}\right].$ (20)
@@
$\mathbb{E}[\sqrt{2}S_{tK+i}|\tilde{Z}_{0:Kt+i-1},Y_{0:Kt+i-1},X_{0}]=0.$
@@
$\beta:=\|K\Gamma_{\tfrac{\alpha}{K}}^{-1}G_{\tfrac{\alpha}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})]\|.$ (21)
@@
$\Sigma:=2\mathbb{E}[S_{tK+i}S_{tK+i}^{\intercal}|\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1}X_{0}].$ (22)
@@
$\sqrt{2}S_{tK+i}$ takes its values in a one-dimensional sub-space almost surely when conditioned on $\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1},X_{0}$ . It is independent of $Z_{tK+i}$ conditioned on $\tilde{Z}_{0:Kt+i-1},Z_{0:Kt+i-1},X_{0}$.
Let us now define the following random probability distributions measurable with respect to the sigma algebra of $\tilde{Z}_{0:Kt+i-1},X_{0}$:
@@
    $Q:=\mathsf{Law}(Y_{Kt+i,2}+\sqrt{2}S_{Kt+i}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$.
@@
    $P:=\mathsf{Law}(\sqrt{I+\Sigma}Z_{Kt+i,2}|\tilde{Z}_{0:Kt+i-1},Z_{0,Kt+i-1},X_{0})$.
@@
$\mathcal{W}_{2}^{2}\left(\bar{P},\bar{Q}\right)\leq\mathbb{E}[\mathcal{W}_{2}^{2}\left(\bar{P},Q\right)|\tilde{Z}_{0:Kt+i-1},X_{0}].$
$\leq 2\mathbb{E}[\mathcal{W}_{2}^{2}\left(\bar{P},P\right)+\mathcal{W}_{2}^{2}\left(P,Q\right)|\tilde{Z}_{0:Kt+i-1},X_{0}].$ (23)
@@
$\leq C\left[\left[\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}\right]\mathbbm{1}(\beta\leq 1)+\frac{\beta^{2}}{K}\mathbbm{1}(\beta>1)\right].$ (25)
@@
$\leq C\left[\tfrac{\beta^{4}}{K^{2}}+\tfrac{\beta^{10}}{K^{3}}+\tfrac{\beta^{6}}{K^{2}}+\tfrac{\beta^{2r}}{K}\right].$ (26)